2.(10 points) Perkins Company produces and sells a single product. The companys income statement for the most recent month is given below: $435,000 Less variable costs: S60,000 Direct labor (variable).. 75,000 45,000 30,000210,000 Variable selling and other expenses.... 225,000 Less fixed expenses: Fixed factory overhead.. Fixed selling and other expenses........ 85,000 185,000 100,000 $40,000 There are no beginning or ending inventories Required a. b. Compute the companys break-even point in units and sales dollars What would the companys monthly net operating income be if sales and total variable costs increased by 30% and total fixed factory overhead dropped by $25,000? What total level of sales (in units) must the company achieve in order to earn a target profit of $145,000? c. d. The company has decided to automate a portion of its operations. The change will reduce direct labor costs per unit by 50 percent, but it will double the costs for fixed factory overhead. Every other cost remains unchanged. Compute the new break-even point in units
Answer #1

a) Break even point in units and sales dollar :

Contribution margin per unit = 225000/15000 = 15

Break even point Units = 185000/15 = 12333 Units

Break even Sales = 12333*29 = $357657

b) Net operating income = (435000*130%)-(210000*130%)-(185000-25000) = 132500

c) Required unit Sales = (185000+145000)/15 = 22000 Units

d) Variable cost per unit = (60000+37500+45000+30000)/15000 = 11.50 per unit

Contribution margin per unit = 29-11.5 = 17.50 per unit

Fixed cost = 200000+85000 = 285000

New break even point = 285000/17.50 = 16286 Units